U.S. Steel is considering a plant expansion to produce austenitic, precipitation hardened, duplex, and martensitic stainless steel round bars that is expected to cost $13 million now and another $10 million 1 year from now. If total operating costs will be $1.6 million per year starting 1 year from now, and the estimated salvage value of the plant is virtually zero, how much must the company make annually in years 1 through 12 to recover its investment plus a return of 16% per year?

The company must make $ million annually in years 1 through 12 to recover its investment plus a return of 16% per year.
Answer #1

Ans.

Present worth (PW) of costs ($ million) = 13 + 10 x P/F(16%,1) + 1.6 x P/A(16%, 12)

= 13 + 10 x 0.86206896551 + 1.6 x 5.19710722221

= 13 + 8.6206896551 + 8.31537155553 = 29.9360612106 or 29.94

Annual income ($ million) = PW / P/A(16%, 12) = 29.94 / 5.19710722221 = 5.76